Pagley Company’s standard labour cost of producing one unit of product DD is 4.35 hours at the rate of $10.35 per hour. During August, 56,320 hours of labour are incurred at a cost of $10.85 per hour to produce 12,800 units of product DD.

Calculate the total labour variance.

Calculate the labour price and quantity variances.

Calculate the labour price and quantity variances, assuming the standard is 4.60 hours of direct labour at $11.00 per hour.
Answer #1

1.

Total labor variance = Standard hour*Standard rate - Actual hours*Actual rate

Standard hours = 12,800*4.35 = 55,680 hours

Total labor variance = 55,680*$10.35 - 56,320*$10.85

Total labor variance = $576,288 - 611,072 = $34,784 Unfavorable

Labour price variance = Actual hours*Standard rate - Actual hours*Actual rate

Labour price variance = 56,320*$10.35 - 56,320*$10.85

Labour price variance = $582,912 - 611,072 = $28,160 Unfavorable

Labour efficiency variance = Standard hour*Standard rate - Actual hours*Standard rate

Labour efficiency variance = $576,288 - 582,912 = $6,624 Unfavorable

2.

Labour price variance = Actual hours*Standard rate - Actual hours*Actual rate

Labour price variance = 56,320*$11 - 56,320*$10.85

Labour price variance = $619,520 - 611,072 = $8,448 Favorable

Labour efficiency variance = Standard hour*Standard rate - Actual hours*Standard rate

Standrad hours = 12,800*4.60 = 58,880 hours

Labour efficiency variance = 58,880*$11 - 56,320*$11

Labour efficiency variance = $647,680 - 619,520 = $28,160 Favorable

Labour efficiency variance =